Determine the point estimate of the population mean and margin of error for the confidence interval. Lower bound is 17, upper bound is 25. The point estimate of the population mean is The margin of error for the confidence interval is
Answer #1

We have given,

Interval = (17,25)

Confidende interval width = Upper bound - Lower bound

= 25-17 = 8

  • Margin of error = rac{Interval , width}{2} = rac{8}{2} = 4
  • Point estimate of pouplation mean is ar{x}

We know that, 100(1-α)% interval is

r t Margin o f error

Lower bound = T-Margin o f error

= Lower bound+ Margin of error = 17 + 4 = 21

Also,

Upper bound = + M argin of error

ar{x} =Upper , bound - Margin , of , error = 25-4 =21

Answe is,

  1. Point estimate of the population mean is 21
  2. Margin of error is 4